Five point charges are enclosed in a cylindrical surface S. If the values of the charges are q1= +3 nC, q2=-2 nC, q3=+2 nC, q4=+4 nC, and q5=-1 nC, find the total flux through S.


1
Expert's answer
2021-04-26T17:13:08-0400
"\Phi=\frac{Q}{\epsilon_0}\\\Phi=\frac{(3-2+2+4-1)10^{-9}}{8.85\cdot10^{-12}}=678\ Vm"